# Black Hole Solutions in Dark Photon Models with Higher Order Corrections

### v1

@@ I Introduction
The nature of dark matter (DM) remains one of the central puzzles in modern particle physics and cosmology. Dark photons are hypothetical spin-1 gauge bosons associated with a hidden U(1) symmetry, proposed in many extensions of the Standard Model. They have attracted increasing interest as candidates for new physics, from dark matter to subtle quantum interference effects Holdom (1986). On the conceptual side, Padavic-Callaghan has recently suggested that the classic double-slit experiment may be reinterpreted in terms of a dark-photon-mediated “phase memory” effect, wherein a hidden photon field carries and preserves the relative phase information of a quantum particle to produce the observed interference pattern Villas-Boas et al. (2025). On the experimental side, searches for dark photon dark matter have yielded new results. An et al. An et al. (2025) report in situ measurements using NASA’s Parker Solar Probe to search for an ambient ultralight dark photon field in the solar corona. While no definitive signal was observed, their data set the strongest constraints to date on dark photons in the $\sim 10^{-9}$ eV mass range (implying an extremely small kinetic mixing, $\chi\sim 10^{-13}$) and approach the sensitivity needed to detect the characteristic resonant conversion “bump” signature of dark photon dark matter. Such developments, bridging quantum laboratory phenomena and astrophysical observations, highlight the timely relevance of investigating dark photon models more deeply on theoretical grounds.

### v2

@@ I Introduction
The nature of dark matter (DM) remains one of the central puzzles in modern particle physics and cosmology. Dark photons are hypothetical spin-1 gauge bosons associated with a hidden U(1) symmetry, proposed in many extensions of the Standard Model. They have attracted increasing interest as candidates for new physics, from dark matter to subtle quantum interference effects Holdom (1986). On the conceptual side, Padavic-Callaghan has recently suggested that the classic double-slit experiment may be reinterpreted in terms of a dark-photon-mediated “phase memory” effect, wherein a hidden photon field carries and preserves the relative phase information of a quantum particle to produce the observed interference pattern Villas-Boas et al. (2025). On the experimental side, searches for dark photon dark matter have yielded new results. An et al. An et al. (2025) report in situ measurements using NASA’s Parker Solar Probe to search for an ambient ultralight dark photon field in the solar corona. While no definitive signal was observed, their data set the strongest constraints to date on dark photons in the $\sim 10^{-9}$ eV mass range (implying an extremely small kinetic mixing, $\chi\sim 10^{-13}$) and approach the sensitivity needed to detect the characteristic resonant conversion “bump” signature of dark photon dark matter. Such developments, bridging quantum laboratory phenomena and astrophysical observations, highlight the timely relevance of investigating dark photon models more deeply on theoretical grounds. Galactic dark matter halos can substantially influence both the generation and propagation of gravitational waves near black holes, underscoring the necessity of incorporating environmental dark matter effects in precise modeling of black hole signals Cardoso et al. (2022). The presence of generic dark matter profiles around black holes alters their apparent image and modifies the characteristics of emitted gravitational waves, highlighting dark matter’s crucial role in interpreting black hole observations Figueiredo et al. (2023). Modeling the central spiky distribution of galactic dark matter as an Einstein cluster reveals how concentrated dark matter cores can affect black hole spacetime properties, emphasizing dark matter’s impact on black hole dynamics Maeda et al. (2025). Treating dark matter as a fundamental vector field within the Einstein cluster framework provides an action principle that links dark matter distributions directly to black hole gravitational behavior, illustrating dark matter’s foundational importance in black hole theory Fernandes and Cardoso (2025). Polar perturbations and energy flux analyses for black holes immersed in generic matter distributions demonstrate that dark matter’s surrounding structure significantly shapes the energy emission and stability of black hole spacetimes Speeney et al. (2024). Investigating a black hole embedded in the Dekel-Zhao dark matter profile illustrates how specific dark matter distributions can tailor the spacetime geometry and potentially observable signatures of black holes Övgün and Pantig (2025).